### v1

# Image Demoireing in RAW and sRGB Domains

Abstract
Moiré patterns frequently appear when capturing screens with smartphones or cameras, potentially compromising image quality. Previous studies suggest that moiré pattern elimination in the RAW domain offers greater efficiency compared to demoiréing in the sRGB domain. Nevertheless, relying solely on raw data for image demoiréing is insufficient in mitigating color cast due to the absence of essential information required for color correction by the Image Signal Processor (ISP). In this paper, we propose perform Image Demoiréing concurrently utilizing both RAW and sRGB data (RRID), which is readily accessible in both smartphones and digital cameras. We develop Skip-Connection-based Demoiréing Module (SCDM) with specific modules embeded in skip-connections for the efficient and effective demoiréing of RAW and sRGB features, respectively. Subsequently, we propose RGB Guided Image Signal Processor (RGISP) to incorporate color information from coarsely demoiréd sRGB features during the ISP stage, assisting the process of color recovery. Extensive experiments demonstrate that our RRID outperforms state-of-the-art approaches by 0.62dB in PSNR and 0.003 in SSIM, exhibiting superior performance both in moiré pattern removal and color cast correction.
1 Introduction
Figure 1: Comparison of image demoiréing in different domains. (a) Image demoiréing in sRGB domain. (b) Image demoiréing in RAW domain. (c) Our proposed RRID, which performs image demoiréing in both RAW domain and sRGB domain
With the increasing prevalence of smartphone cameras, it is common to record on-screen content using smartphone cameras for convenience. However, this type of photography often results in the presence of moiré patterns in the captured images. The appearance of moiré patterns is caused by spatial frequency aliasing between the camera’s color filter array (CFA) and the screen’s LCD subpixel, resulting in an unsatisfactory visual experience.
Moiré patterns present challenges to be eliminated due to their different scales, indistinct shapes, diverse colors, and varying frequencies. Nowadays, numerous learning-based image demoiréing approaches [20, 2, 6, 7, 15, 32, 16, 22, 26, 17, 21] have been proposed. Among the various methods employed, a commonly used approach is utilizing a multi-scale architecture to remove various sizes of moiré patterns [20, 2, 6, 26, 17]. Besides, some works advocate the leverage of frequency components and delicately design frequency-based demoiréing networks [7, 15, 32, 21].
The aforementioned methods are developed to remove moiré patterns in the standard RGB (sRGB) domain, as depicted in Fig. 1(a). While this scheme seems to be the most direct approach, its effectiveness is limited in removing moiré patterns from complex scenes. The reason for this limitation probably lies in the deterioration of moiré patterns in the RAW domain, caused by the nonlinear operation during the Image Signal Processor (ISP) process, such as demosaicing. Prior researches consequently suggest that eliminating moiré patterns in the RAW domain is more efficient than performing demosaicing in the sRGB domain [27, 29]. Due to the easy accessibility of RAW data in both smartphones and digital cameras, utilizing it for image demoiréing can be both feasible and advantageous. RDNet [27] is the pioneering study that investigates image demoiréing in the RAW domain. The researchers explore the demoiréing of RAW images using an encoder-decoder and class-specific learning approach, followed by the utilization of a pre-trained ISP for the RAW to sRGB domain conversion process, as depicted in Fig. 1(b). Nevertheless, relying exclusively on RAW data can lead to color cast in the restored images due to the absence of essential information in the ISP.
In order to correct the color cast and make the best use of RAW and sRGB information for moiré pattern removal, we propose image demoiréing in both the RAW and sRGB domains (RRID). Firstly, we advocate that introducing RAW data is advantageous moiré removal due to the following reasons: (1) RAW pixels provide more information compared to sRGB pixels since they are typically 12 or 14 bits. (2) Moiré patterns in the RAW domain are less apparent because they are not further affected by nonlinear operations in the ISP. In addition, we further incorporate sRGB data as the color transformation guidance in the ISP, as demonstrated in Fig. 1(c), facilitating color deviations correction. In RRID, we adopt a multi-scale architecture in the Skip-Connection-based Demoiréing Module (SCDM), to effectively eliminate moiré patterns of different sizes. In order to efficiently perform demoiréing on both RAW data and sRGB features, distinct modules are integrated into skip-connections within SCDM, as opposed to brutally injecting them into the multi-scale layers. Specifically, for RAW data demoiréing, we introduce the Gated Feedback Module (GFM) into the skip connection, enabling adaptive differentiation between texture details and moiré patterns through feature gating. Moreover, we design the Frequency Selective Module (FSM) for moiréd sRGB features, leveraging a learnable band stop filter to mitigate moiré patterns in the frequency domain. Subsequent to the pre-demoiréing operations, we develop the sRGB Guided Image Signal Processor (RGISP) to integrate color information from the coarsely demoiréd sRGB features with the coarsely demoiréd RAW features during the ISP stage, thereby facilitating the color recovery process. Ultimately, for image reconstruction, we utilize multiple Residual Swin Transformer Blocks (RSTBs) [13] and convolutions to accomplish global tone mapping and detail refinement.
In summary, our contributions are listed as follows:
  • $\bullet$
    We propose to exploit image demoiréing in both RAW and sRGB domains. The incorporation of RAW data enhances the moiré pattern removal process, while the inclusion of sRGB data reduces the difficulty of RAW-to-sRGB inversion and aids in color restoration of images.
  • $\bullet$
    We develop SCDM with specifically designed modules embeded in the skip-connections to perform efficient and effective demoiréing on the RAW and sRGB features.
  • $\bullet$
    RGISP is designed for RAW-to-sRGB conversion, incorporating color information from coarsely demoiréd sRGB feature with demoiréd RAW features during the ISP stage, facilitating the process of color deviation correction.
  • $\bullet$
    Extensive experiments are conducted on the RAW demoiréing dataset. Our RRID surpasses state-of-the-arts methods in both qualitative and quantitative evaluations.
2 Related Works
2.1 Image and Video Demoiréing
To mitigate moiré patterns resulting from frequency aliasing, numerous learning-based image demoiréing methods [2, 7, 32, 16, 22, 17, 21] and video demoiréing methods [3, 18, 25, 29] have been introduced. DMCNN [20] introduces the first real-world dataset for image demoiréing along with a multi-resolution convolutional neural network. MopNet [6] devises a multi-scale aggregated network guided by edges and attentive to pattern attributes for moiré pattern removal. WDNet [15] employs wavelet transformation to decompose images with moiré patterns into separate frequency bands and establishes a dual-branch network to restore both close-range and far-range information. FHDMi [7] formulates a two-stage approach for simultaneous removal of substantial moiré patterns while preserving image details. ESDNet [26] explores a lightweight model aimed at ultra-high-definition image demoiréing. VDMoiré [3] conducts the first video demoiréing dataset and presents a baseline video demoiréing model that utilizes three consecutive frames. The aforementioned methods all operate in the sRGB domain. However, it has been observed that moiré patterns are less prominent in the RAW domain, which aids in the process of moiré pattern removal. Therefore, some researchers advocate performing image or video demoiréing in the RAW domain. RDNet [27] introduces the first image demoiréing dataset with Raw data and conducts the demoiréing process in the RAW domain. RawVDemoiré [29] proposes a temporal alignment method for Raw video demoiréing. To effectively remove moiré patterns in the RAW domain, the model must possess the simultaneous capabilities of demoiréing and domain conversion. However, previous methods encounter challenges in color recovery due to the limited availability of relevant information in the ISP. Consequently, we propose an image demoiréing network that utilizes both sRGB and Raw data, explicitly addressing the aforementioned issues.
Figure 2: The overview and detailed structures of our proposed RRID for image demoiréing.
2.2 Learning-Based RAW Image Processing
Raw pixels offer the potential for leveraging additional information due to their broader bit width, inherently containing a greater wealth of data. Early research in Raw image processing primarily centers on integrating demosaicking and denoising techniques [10, 5, 11, 14]. RTF [10] demonstrates a machine learning approach to the demosaicing problem, providing versatility by being applicable to various tiled CFA patterns. SGNet [14] introduces a self-guidance network that incorporates density-map guidance and green-channel guidance for the combined tasks of demosaicing and denoising. Recently, numerous low-level vision tasks endorse the utilization of Raw data, encompassing low-light enhancement [8, 4, 9], super-resolution [24, 28], and reflection removal [12, 19]. The majority of the previously mentioned Raw image processing methods only take Raw images as input and generate restored sRGB images as output, indicating that the entire model is responsible for both the restoration and ISP tasks. CR3Net [19] presents a cascaded network that harnesses both the RGB data and their corresponding Raw versions for image reflection removal. Within CR3Net, a module is formulated to emulate point-wise mappings within the ISP, translating Raw features into the sRGB domain. To further mitigate the color casting problem, our proposed method incorporates guidance from sRGB information with the Raw information within the ISP process.
3 Methodology
The purpose of image demoiréing is to restore the original clean image by removing moiré patterns and correcting color deviations from an contaminated image. Originally caused by frequency interference between the camera’s CFA and the screen’s LCD subpixel, moiré patterns are further deteriorate by the non-linear processes in ISP. Therefore, introducing raw data for image demoiréing may offer greater efficiency. However, relying solely on raw data for image demoiréing is insufficient in mitigating color cast due to the absence of essential information required for color correction by the Image Signal Processor (ISP). Subsequently, we propose to concurrently utilize both RAW and sRGB data, which is readily accessible in both smartphones and digital cameras, for image demoiréing.
Fig. 2 presents an overview of our proposed RRID for image demoiréing. The input of the RGB stage is an sRGB image $\mathbf{I}_{rgb}\in\mathbb{R}^{H\times W\times 3}$, where $H\times W$ represents the spatial resolution. Given an input RAW image, it is first packed into the 4-channel RGGB format $\mathbf{I}_{raw}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 4}$, which serves as the input for the RAW stage. RRID is designed to generate both the final output sRGB image ($\mathbf{\hat{Y}}_{rgb}$) and a RAW image ($\mathbf{\hat{Y}}_{raw}$) for supervision.
Concretely, shallow features $\mathbf{F}_{raw}$ and $\mathbf{F}_{rgb}$ are initially extracted using a convolutional layer and the Dilated Channel Attention Block (DCAB). It is worth noting that a down-sampling layer is applied to the sRGB data to ensure the uniformity of feature shapes. Then, we apply SCDM to acquire pre-demoiréd features $\mathbf{D}_{raw}$ and $\mathbf{D}_{rgb}$ in RAW stage and RGB stage, respectively. GFM is introduced in the skip-connection of SCDM for raw features, enabling adaptive differentiation between texture details and moiré patterns through feature gating. Similarly, FSM is utilized to to mitigate moiré patterns in the frequency domain. More details will be given in Sec. 3.1. Subsequently, RGISP is designed to realize a raw-to-rgb conversion, acquiring $\mathbf{D}_{out}$. RGISP incorporates color information from coarsely demoiréd sRGB features $\mathbf{D}_{rgb}$ in a cross-attention mechanism, which will be introduced in Sec. 3.2. For the final image reconstruction, 4 RSTBs [13] are utilized for their advantages in building long-range dependencies for global tone mapping and detail refinement.
3.1 Design of SCDM
We propose SCDM, a multi-scale architecture designed to remove moiré patterns from shallow RAW features $F_{raw}$ and shallow sRGB features $F_{rgb}$, as shown in Fig. 2(a). Unlike the traditional vanilla U-Net, the SCDM is constructed upon DCAB, facilitating the encoding and decoding of informative features within the respective layers. The features are downsampled twice in the spatial dimension during the encoding process and then upsampled to the original resolution of the input features during decoding. To emphasize the contextual information, we enhance the learned features from the corresponding layers of the encoder with additional modules at skip connections, that is GFM for RAW and FSM for sRGB. The introduction of carefully designed demoiréing modules in skip-connections is also intended to enhance model efficiency. Then, the pre-demoiréd features $\mathbf{D}_{raw}$ and $\mathbf{D}_{rgb}$ are acquired.
DCAB. As the essential building block of encoder-decoders for realizing RAW demoiréing and the RGB demoiréing, the detailed structure of DCABs is shown in Fig. 2(b). DCAB is designed using a series of dilated convolution layers followed by ReLU activation and the channel attention mechanism. The receptive field at each block is expanded with the dilated convolution kernels, facilitating the demoiréing operations. Furthermore, by incorporating the channel attention mechanism [31], DCAB adaptively rescales features according to interdependencies among channels. Also, shortcuts are introduced to establish connections between the input and output of DCAB, enabling the block to learn the residual features.
GFM. Embeded in the skip-connection of SCDM for raw features, GFM is introduced to enable adaptive differentiation between texture details and moiré patterns through feature gating. For efficiency, we employ point-wise and depth-wise convolutions to aggregate channel and local content information, respectively. Then, the features are chunked along the channel dimension to $\mathbf{F}_{gate}$ and $\mathbf{F}_{content}$. A point-wise multiplication is applied on $\mathbf{F}_{content}$ and $\mathbf{F}_{gate}$ after a $\mathrm{GELU}$ activation. During feature gating, we expect that the original image content is adaptively selected and merged along both spatial and channel dimensions. The detailed structure of the GFM is shown in Fig. 2(c). The operations can be formulated as:
$\mathbf{F}_{gate},\mathbf{F}_{content}=\mathrm{DConv}(\mathrm{PConv}(\mathbf{F}_{GFM\_in})),$ (1)
$\mathbf{F}_{GFM\_out}=\mathrm{PConv}(\mathbf{F}_{content}\odot\mathrm{GELU}(\mathbf{F}_{gate})+\mathbf{F}_{GFM\_in})$ (2)
where $\mathrm{DConv}$ and $\mathrm{PConv}$ represent depth-wise convolution and point-wise convolution, respectively. $\odot$ denotes the element-wise multiplication.
Table 1: Quantitative comparison with the state-of-the-art demoiréing approaches and raw image restoration methods on TMM22 dataset [27] in terms of average PSNR, SSIM, LPIPS and computational complexity. The best results are highlighted with bold. The second-best results are highlighted with underline.
Index Methods
DMCNN WDNet ESDNet RDNet DNF RawVDmoiré CR3Net RRID
[20] [15] [26] [27] [9] [29] [19] (Ours)
# Input type sRGB sRGB sRGB RAW RAW RAW sRGB+RAW sRGB+RAW
[…]
Params (M) 1.55 3.36 5.93 6.04 2.80 5.33 2.68 2.46
TFLOPs 0.095 0.882 4.478 0.161 0.013 0.341 0.221 0.093
Inference time 0.052 0.284 0.115 1.094 0.070 0.182 0.058 0.089
FSM. For sRGB demoiréing, we leverage a learnable band stop filter [32] to mitigate moiré patterns in the frequency domain. Considering the formation of moiré patterns is closely related to the camera imaging process, especially the frequency of the color filter array, demoiréing in the frequency domain becomes a favorable approach. However, the frequency spectrum of the moiré pattern is often entangled with that of the original contents, thereby increasing the difficulty of disentanglement. Therefore, we propose to perform the disentanglement within small patches, where the moiré patterns tends to be more pronounced in a specific range of frequency bands, thereby facilitating the removal of moiré patterns. Inspired by Zheng et al. [32], we utilize a band stop filter to amplify certain frequencies and suppress others using Block-DCT. However, obtaining the frequency spectrum prior to model the moiré texture is challenging, and determining a frequency prior for each moiré image patch is time-consuming. To address this issue, a set of convolution layers are employed as the learnable band stop filter $\mathrm{B}(\cdot)$ to attenuate the specific frequencies of moiré patterns while preserving the original image contents. The process of the FSM can be represented as:
$\mathbf{F}_{FSM\_out}=\mathrm{IDCT}(\mathrm{B}(\mathrm{DCT}(\mathbf{F}_{FSM\_in}))),$ (3)
where $\mathrm{DCT}$ and $\mathrm{IDCT}$ denote the process of Block-DCT and the corresponding inverse process.
3.2 Design of RGISP
RGISP is proposed to transform pre-demoiréd RAW features $\mathbf{D}_{raw}$ into sRGB domain $\mathbf{D}_{out}$, with the assistance of pre-demoiréd sRGB features $\mathbf{D}_{rgb}$. Matrix transformation is commonly employed in the traditional ISP pipelines [1].It enables the enhancement or conversion of image colors to another color space through a channel-wise matrix transformation, which is facilitated by globally shared settings like environmental illumination and color space specifications [9]. Building upon this principle, we introduce matrix transformation as a means to perform color transformation, as illustrated in Fig. 3. The design of this matrix transformation is informed by recent advancements in transposed cross-attention.
Figure 3: The structure of RGISP.
Given $\mathbf{D}_{raw}\in\mathbb{R}^{C\times\frac{H}{2}\times\frac{W}{2}}$, the vectors of the query $\mathbf{Q}\in\mathbb{R}^{C\times\frac{H}{2}\frac{W}{2}}$, key $\mathbf{K}\in\mathbb{R}^{\frac{H}{2}\frac{W}{2}\times C}$ are generated through the projection with a 1 $\times$ 1 convolutional layer $\mathrm{Conv}$ and a flatten operation $\mathrm{Flatten}$, where $C$ is the intermediate channel number. Similarily, value $\mathbf{V}\in\mathbb{R}^{C\times\frac{H}{2}\frac{W}{2}}$ can be acquired by $\mathbf{D}_{rgb}\in\mathbb{R}^{C\times\frac{H}{2}\times\frac{W}{2}}$ as:
$\mathbf{Q},\mathbf{K}=\mathrm{Flattern}(\mathrm{Conv}(\mathbf{D}_{raw})),$ (4)
$\mathbf{V}=\mathrm{Flatten}(\mathrm{Conv}(\mathbf{D}_{rgb})).$ (5)
Then, the transformation matrix $\mathbf{M}\in\mathbb{R}^{C\times C}$ is obtained by matrix multiplication. The procedure can be formulated as:
$\mathbf{M}=\mathrm{Softmax}(\mathbf{Q}\cdot\mathbf{K}^{T}/\lambda),$ (6)
where a scaling coefficient $\lambda$ is applied for numerical stability. Subsequently, the vector $\mathbf{V}$ is transformed by the matrix $\mathbf{M}$, performing color space conversion in feature-level. The output feature after color transformation can be obtained by $\mathbf{D}_{out}=\mathbf{M}\cdot\mathbf{V}$. As a complement to the global matrix transformation, we use a depth-wise convolution and a point-wise convolution to refine the local details.
3.3 Loss Function
To effectively remove the moiré patterns and adjust the color deviations, we introduce supervision on different domains, clean RAW and clean sRGB. The overall training objective can be expressed as:
$\small\mathcal{L}=\alpha\|\hat{\mathbf{Y}}_{raw}-\mathbf{Y}_{raw}\|_{1}+\|\hat{\mathbf{Y}}_{rgb}-\mathbf{Y}_{rgb}\|_{1},$ (7)
where ${\mathbf{Y}}_{raw}$ and ${\mathbf{Y}}_{rgb}$ are the ground truth RAW image and sRGB image. Only $L_{1}$ loss is employed for both RAW and sRGB supervision. Since our goal is to restore the image in the sRGB domain, we empirically set $\alpha$=0.5.
4 Experiments
4.1 Experimental Setup
Dataset. The experiments are conducted on the RAW image demoiréing dataset published in TMM22 [27], which contains recaptured scenes of natural images, documents, and webpages. The moiré images are captured using four different smartphone cameras and three display screens of varying sizes. In TMM22, the dataset consists of 540 RAW and sRGB image pairs with ground-truth for training, and 408 pairs for testing. To facilitate the training and comparison process, patches of size 256$\times$256 and 512$\times$512 are cropped respectively for the training and testing sets.
Training Details. The proposed RRID utilizes both RAW and sRGB images as inputs to generate the restored image in the sRGB domain. The model is trained using the AdamW optimizer with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999. A multistep learning rate schedule is employed, with the learning rate initialized at 2×$\times 10^{-4}$. The RRID model is trained for 500 epochs using a batch size of 80 on 4 NVIDIA RTX 3090 GPUs.
4.2 Quantitative Results
We compare our approach with several image demoiréing methods in sRGB domain, including DMCNN [20], WDNet [15], ESDNet [26], a RAW image demoiréing method RDNet [27]. We further add comparison with a low-light enhancement method DNF [9] with RAW input, a video demoiréing method RawVDmoiré [29] with RAW input, and a deflection removal method CR3Net [19] with paired RAW and sRGB data. To evaluate the performance of our proposed RRID, we adopt the following three standard metrics to assess pixel-wise accuracy and the perceptual quality: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [23], and Learned Perceptual Image Patch Similarity (LPIPS) [30]. Additionally, the number of parameters, Floating-Point Operations per Second (FLOPs), and inference time are introduced to compare the model complexity. To ensure a fair comparison, we adjust the inputs of certain models to align with the dataset requirements. Subsequently, all models are fine-tuned using the default parameter settings as provided in their paper on the TMM22 dataset. The superior results are reported between the pre-trained and fine-tuned models.
Table. 1 presents the quantitative and the computational complexity comparisons. Our proposed RRID outperforms the second-best method, RawVDmoiré, 0.62dB in PSNR and 0.003 in SSIM. As for LPIPS, our methods accomplishes an acceptable result at 0.079, surpassing most of the comparison methods. Even though our method utilizes both RAW data and sRGB data as inputs, our approach can still achieve a satisfactory demoiréing effect with a inference time of merely 0.089s. Compared to previous methods, our RRID demonstrates superiority by employing relatively fewer parameters and FLOPs, further demonstrating the superiority of our approach.
Table 2: Quantitative comparison on FHDMi dataset [7].
[…]
DMCNN [20] 21.54 0.773
MopNet [6] 22.76 0.976
MBCNN [32] 22.31 0.810
[…]
ESDNet [26] 24.50 0.835
Ours 24.39 0.830
[…]
4.3 Qualitative Results
We present visual comparisons between our approach and the existing methods in Fig. 4. The results demonstrate the advantage of RRID in effectively removing moiré artifacts and correcting color deviations. For the first scene demonstrating a recaptured flower with moiré patterns, previous methods tend to exhibit residual moiré patterns in their restored images. While ESDNet successfully eliminates a significant portion of these patterns, it struggles to accurately restore the original colors of the image. Apart from natural scenes, we also present a case of webpages contaminated with moiré patterns. Through the utilization of both RAW and sRGB data, our approach adeptly eliminates moiré patterns while preserving fine image details, even in scenarios characterized by severe color deviations. This ensures a more accurate restoration of the original colors and results in enhanced visual effects. Additional qualitative comparisons can be found in the supplementary file.
Figure 4: Qualitative comparison on raw image demoiréing TMM22 dataset [27].
4.4 Ablation Study
1) Ablation study on the model architecture and inputs.
Table 3: Baseline models comparison for the evaluation of the architecture and the inputs.
[…]
$B_{1}$ (w/o RAW in and RAW branch) 25.79 0.915
$B_{2}$ (w/o sRGB in and sRGB branch) 27.24 0.929
$B_{3}$ (w/o RAW in) 26.51 0.923
$B_{4}$ (w/o RGISP) 27.38 0.932
$B_{5}$ (w/o Reconstruction) 26.93 0.926
$B_{6}$ (full version) 27.88 0.938
In Table 3, we investigate the architecture and inputs of RRID and validate the importance of different individual components in the whole RRID. We conduct this ablation study by comparing the proposed RRID and the following variants of RRID:
  • $\bullet$
    $B_{1}$. Both the RAW input and the RAW demoiréing branch are removed.
  • $\bullet$
    $B_{2}$. Both the sRGB input and sRGB demoiréing branch are removed.
  • $\bullet$
    $B_{3}$. The sRGB input is substituted with RAW input. The input channel number is adjusted, while the rest of the architecture is unchanged.
  • $\bullet$
    $B_{4}$. We replace RGISP with a two convolutional layers after a concatenation of the RAW and sRGB features.
  • $\bullet$
    $B_{5}$. Reconstruction module is replaced with two convolutional layers.
  • $\bullet$
    $B_{6}$. The full model of RRID with both RAW and sRGB images as model input.
Figure 5: Visualization of the ablation study for architecture and inputs comparison.
Our proposed RRID leverages both RAW data and sRGB as input, employing distinct demoiréing modules for each data type to extract pre-demoiréd features. Removing the RAW input and its corresponding raw demoiréing module ($B_{1}$) results in a drastic decrease in PSNR of approximately 2dB. Likewise, eliminating the sRGB input and its corresponding RGB demoiréing module ($B_{2}$) leads to a PSNR of only 27.24dB. In addition to the decline in numerical metrics, we can also assess the visual performance from Fig. 5. When relying only on sRGB data and its corresponding demoiréing modules, moiré patterns are less effectively removed, as depicted in Fig.5(b). Conversely, Fig.5(c) demonstrates that using only raw data as input results in images with more pronounced color deviations. To further validate the significance of RAW input for image demoiréing, we compare ($B_{3}$) to ($B_{1}$) by solely removing the RAW input data while retaining the remaining model structures. Although ($B_{3}$) demonstrates some improvement in numerical metrics compared to ($B_{1}$), its performance remains unsatisfactory due to the lack of the additional rich original information provided by the RAW input and the fact that the original raw demoiréing module is specifically designed for RAW data. Additionally, we conducted further comparisons by removing the originally designed RGISP ($B_{4}$) and Reconstruction modules ($B_{5}$), resulting in inferior numerical metrics and performance in Fig. 5 compared to the complete RRID model ($B_{6}$).
2) Ablation study on SCDM.
Table 4: Ablation study on SCDM.
[…]
$S_{1}$ (w/o GFM & FSM) 27.00 0.926
$S_{2}$ (w/o GFM) 27.17 0.927
$S_{3}$ (w/o FSM) 27.39 0.930
$S_{4}$ (all GFM) 27.50 0.933
$S_{5}$ (all FSM) 27.32 0.930
$S_{6}$ (w/o DCAB) 26.55 0.923
$S_{7}$ (SCDM) 27.88 0.938
SCDM is designed to perform efficient demoiréing on both raw data and sRGB data, with distinct modules incorporated into skip-connections.More specifically, for raw feature demoiréing, we introduce GFM into the skip connection, enabling adaptive differentiation between texture details and moiré patterns through feature gating. Similarily, FSM is embeded in SCDM for sRGB moiréd features, leveraging a learnable band stop filter to mitigate moiré patterns in the frequency domain.
Figure 6: Visualization of the ablation study for SCDM.
Table 4 demonstrates the effectiveness and indispensability of the meticulously designed GFM and FSM integrated within the skip-connections. Removing the demoiréing modules on the skip-connections ($S_{1}$-$S_{3}$) individually results in varying degrees of performance decline for the model. Fig.6(b) demonstrates that the model’s ability to remove moiré patterns decreases upon removing the GFM and FSM modules. Furthermore, to assess the specificity of GFM and FSM, we substitute the original modules in $S_{4}$ and $S_{5}$, respectively. Despite no reduction in the number of parameters, the PSNR decreased by approximately 0.3-0.5dB. Since DCAB is a crucial component module in SCDM, we substitute DCAB with two convolutional layers to confirm its necessity. Fig. 6(c) showcases more residual moiré patterns and significant color deviation compared to our complete SCDM presented in Fig.6(d).
3) Ablation study on RGISP.
Table 5: Ablation study on RGISP.
[…]
$R_{1}$ (self-atten([raw, rgb])) 27.54 0.933
$R_{2}$ ([self-atten(raw), rgb]) 27.42 0.931
$R_{3}$ (RRM) 27.21 0.929
$R_{4}$ (RGISP) 27.88 0.938
Figure 7: Visualization of the ablation study for RGISP.
RGISP is designed to integrate color information from the coarsely demoiréd sRGB features during the ISP stage, facilitating the color recovery process. The process of the RGISP can be recognized as a cross-attention mechanism. To assess the efficacy of the RGISP design, we compare the following structures in Table 5. For $R_{1}$, we concatenate the pre-demoiréd RAW features with the sRGB features and subsequently apply a self-attention mechanism. In $R_{2}$, the pre-demoiréd RAW features undergo a self-attention mechanism and are then concatenated with the pre-demoiréd sRGB features. Table 5 shows that our proposed RGISP surpasses the performance of self-attention mechanisms ($R_{1}$, $R_{2}$), as well as other raw-to-rgb conversion methods RRM [19] ($R_{3}$), based on numerical metrics. Fig. 7 depicts a webpage with globally distributed moiré patterns. $R_{1}$-$R_{4}$ are all capable of effectively removing the moiré patterns, but $R_{4}$ exhibits the best performance in correcting color deviation. With the introduction of sRGB pre-demoiréd features, RGISP is able to restore the original colors of the image effectively.
5 Conclusion
In this paper, we propose to concurrently utilize both RAW and sRGB data for image demoireing (RRID). Firstly, we introduce SCDM with dedicated modules embedded in skip-connections to handle the demoireing of RAW and sRGB features, respectively. Additionally, we present RGISP, which integrates color information from coarsely demoired sRGB features into the ISP stage to aid in the color recovery process. Extensive experiments demonstrate the superior performance of our RRID method in both moire pattern removal and color cast correction. Our approach outperforms state-of-the-art methods by 0.62dB in PSNR and 0.003 in SSIM.
Supplementary Material
6 Additional Ablation Study for SCDM
In Table 6, we validate the effectiveness of embedding the demoiréing modules in the skip-connections. We conduct this ablation study by comparing the proposed SCDM and the following cases:
  • $\bullet$
    $A_{1}$. In the Raw branch, DCAB is replaced with GFM, while in the RGB branch, DCAB is replaced with FSM. The demoiréing modules embedded in the skip-connections are reserved for both branches.
  • $\bullet$
    $A_{2}$. In the Raw branch, DCAB is replaced with GFM, while in the RGB branch, DCAB is replaced with FSM. The demoiréing modules embedded in skip-connection are removed for both branches.
  • $\bullet$
    $A_{3}$. RRID with the original SCDM design.
From $A_{1}$, it can be observed that replacing DCAB with GFM and FSM results in a noticeable decline in both PSNR and SSIM. A more significant concern arises from the substantial increase in inference time, primarily caused by the block-DCT utilized in FSM. This observation underscores the crucial role of DCAB, a pivotal block in SCDM, in encoding and decoding informative features. Moreover, even when we eliminate the demoiréing modules on the skip-connections in $A_{1}$, the inference time remains relatively long for $A_{2}$. This finding emphasizes the effectiveness and efficiency of incorporating demoiréing modules within skip-connections.
Table 6: Ablation Study for SCDM.
[…]
$A_{1}$ 27.11 0.927 4.589
$A_{2}$ 26.68 0.923 3.909
$A_{3}$ (RRID) 27.88 0.938 0.089
7 Limitation
While we propose employing both RAW and sRGB images for moiré pattern removal, and design RGISP to integrate color information from coarsely demoiréd sRGB features with demoiréd RAW features during the ISP stage, achieving flawless color restoration in scenes with significant color deviations remains a challenging task.
Figure 8: Failure cases in correcting color deviations.
Fig. 8 demonstrates two scenes where the input images containing moiré patterns display noticeable color discrepancies compared to the ground truth images. While our proposed method effectively removes moiré patterns and partially corrects the color, it falls short of fully mapping the colors to those present in the ground truth. Currently, there is only one dataset, TMM22 dataset [27], available for image demoiréing with RAW-RGB pairs. However, the images in the training set of TMM22 are limited to a size of 256$\times$256, which may constrain the model’s capacity to learn global color mapping. To overcome this limitation, one possible solution is to collect a RAW-RGB paired image demoiréing dataset with a higher resolution, which will be our future work.
8 More Qualitative Comparison on TMM22 Dataset
More visual comparisons with state-of-the-arts methods are presented in Figs. 9-10. The results demonstrate the superiority of our proposed RRID in demoiréing and restoring color. For instance, in the first scene depicted in Fig. 9, the input image suffers from both severe moiré pattern contamination and noticeable color deviation. Previous methods struggle to remove moiré patterns and encounter difficulties in accurately restoring the petal color to pink. In contrast, our approach successfully eliminates moiré patterns and effectively corrects color deviations. The last scene illustrated in Fig. 10 presents a moiréd document. Given the relatively slight moiré contamination in this scene, the majority of methods successfully remove the moiré patterns. Notably, the colored area at the top of the image deserves attention as most methods incorrectly correct it as blue. However, the ground truth image reveals that this area is actually cyan. In contrast, our method excels at faithfully restoring the image’s color.
Figure 9: Qualitative comparison on raw image demoiréing TMM22 dataset [27].
Figure 10: Qualitative comparison on raw image demoiréing TMM22 dataset [27].